OpenStudy (anonymous):

The table below shows the cumulative distance and the number of hours Ryan drove on vacation. Hours Driving Distance Traveled (miles) 0 0 3 168 4 224 7 392 10 560 Based on the information in the table, at what speed did he drive (miles per hour)? 46 miles per hour 56 miles per hour 63 miles per hour 39 miles per hour

OpenStudy (anonymous):

@kohai @MARC_

OpenStudy (kohai):

Same thing, divide the miles by the hour, just like the last one

OpenStudy (anonymous):

Okay. thankies ~.~ @kohai but i still dont see how i would divide them?..

OpenStudy (kohai):

Okay, I can explain

OpenStudy (kohai):

When you see things like "miles per hour" or "price per pound," it means: \[\frac{ miles }{ hour }\] \[\frac{ price }{ pound }\]

OpenStudy (kohai):

So, let's look at the first one on your table

OpenStudy (kohai):

560 miles in 10 hours \[\frac{ 560 }{ 10 }\] 56 miles per hour

OpenStudy (kohai):

See how that works?

OpenStudy (anonymous):

Yeah. But basically you just removed the 0..? @kohai sorry i think my brain turned off

OpenStudy (kohai):

Yeah but that's just because they're tens. We can look at another

OpenStudy (anonymous):

Okay. So the answer would be 56 though??

OpenStudy (kohai):

\[\frac{ miles }{ hour }\] \[\frac{ 392 miles }{ 7 hours }\]

OpenStudy (kohai):

Yeah

OpenStudy (anonymous):

Okay. @kohai